An ice cube of mass 65 g is taken from a freezer at -15


oC and dropped into 125 cm3 of water at


85

oC. What will be the final temperature of the mixture? The specific heat of ice is 0.5 cal/g.oC

and latent heat of fusion of ice is 80 cal/g.


1
Expert's answer
2022-02-19T06:55:14-0500

Initial temperature, T1= -15°c

"Q= m\u00d7 c\u00d7 dT"

80= 190× 0.5× ( T2-T1)

80= 190× 0.5( T2+15)

80= 190× 0.5T2+ 7.5

0.421= 0.5T2+ 7.5

7.921= 0.5T2

T2= 15.8°c